OpenStudy (clarence):

My friend got 1/15 whereas I got 15/16 for the following series:

OpenStudy (clarence):

\[\sum_{n=1}^{\infty} \frac{ 1^{n} }{ (-15)^{n-1} }\]

OpenStudy (anonymous):

it can't really be \(1^n\) can it?

OpenStudy (clarence):

It unfortunately is :/

OpenStudy (anonymous):

lol ok you are right your friend is wrong

OpenStudy (clarence):

Just a guess? :P

OpenStudy (anonymous):

no not a guess, you can do it in your head

OpenStudy (clarence):

You're obviously way more educated than I am then

OpenStudy (anonymous):

whats is \(1-(-\frac{1}{15})\)?

OpenStudy (anonymous):

aka \[1+\frac{1}{15}\]?

OpenStudy (clarence):

16/15?

OpenStudy (anonymous):

yes of course and the reciprocal of \(\frac{16}{15}\) is?

OpenStudy (clarence):

15/16?

OpenStudy (anonymous):

no not 15/16? but rather 15/16 !!! how hard was that?

OpenStudy (clarence):

Fair enough! I guess I'm just so used to the teacher saying to do questions a certain way that I didn't think about that, thanks!

OpenStudy (anonymous):

i am not that educated, this is easy is all for example \[\sum\left(\frac{2}{3}\right)^n\] \[1-\frac{2}{3}=\frac{1}{3}\] the reciprocal of \(\frac{1}{3}\) is \(3\) finished math teaches like to make things harder than they actually are